HFI operations timeline

From Planck PLA 2015 Wiki
Jump to: navigation, search

The following list gives some notable events and activities that are known to have affected or might have affected the HFI(Planck) High Frequency Instrument data. The time information is the start time of the event in On Board Time. The label contains some keywords:

  • dtcp, (for Daily Tele-Communication Period) refers to an extended pass;
  • HFI(Planck) High Frequency Instrument, refers to HFI(Planck) High Frequency Instrument instrument operations (when this is during the commissioning of the CPVCalibration and Performance Verification phases, labels are usually followed by the activity label);
  • period, refers to the different phases of the mission;
  • sat, refers to operations on the satellite SerVice Module that might have an impact on the HFI(Planck) High Frequency Instrument data;
  • slew, refers to thruster activity;
  • solar flare, refers to all solar events as detected by a thermal fluctuation on the HFI(Planck) High Frequency Instrument bolometer plate;
  • test, refers to End Of Life tests executed on the HFI(Planck) High Frequency Instrument;
  • therm, refers to satellite events that might have a thermal impact on the HFI(Planck) High Frequency Instrument.


14/05/2009 13:12:00	EPC Vulcain Engine ignition
14/05/2009 16:58:36	HFI: LCLs nominal switch on
14/05/2009 17:30:59	HFI: go from startup to standby 
14/05/2009 17:36:08	HFI: choose DPU FDIR mode Desynchro
14/05/2009 17:39:39	HFI: open dilution cooler valves
14/05/2009 18:48:52	HFI: setup REU Ns (40 Hz) and Fdiv (124) parameters
14/05/2009 18:52:33	HFI: set 4-K cooler internal frequency at 40Hz
14/05/2009 19:05:45	HFI: activate the 6 nominal cernox thermometers
14/05/2009 19:08:51	HFI: synchronise the 4-K cooler clock (External)
14/05/2009 19:23:47	HFI: set the 4-K cooler amplitude demand at 2000 microns
14/05/2009 19:28:42	HFI: power on the nominal 4-K getter heater
14/05/2009 19:35:31	HFI: start the 4-K cooler compressors
15/05/2009 12:00:00	therm: stop the Focal Plane Unit 250K heating
15/05/2009 19:17:06	slew: OCM LEOP (MAN-01a)
16/05/2009 12:06:55	HFI: switch on all REU analog belts
16/05/2009 12:17:23	HFI: REU FPGA reset
16/05/2009 12:26:40	HFI: start the 4-K cooler compressors
16/05/2009 12:31:16	HFI: set the JFET PID heaters at 110K
16/05/2009 12:49:26	HFI: set the channel compression parameters
16/05/2009 12:54:37	HFI: raw signal cycle on channels
16/05/2009 14:20:15	HFI: science data production starts
16/05/2009 14:48:30	HFI: channel setup [all channels]
16/05/2009 15:41:23	HFI: VI measurement [bolos, thermos, resistor]
16/05/2009 16:00:00	HFI: test end of slew
16/05/2009 17:51:43	HFI: set Cernox thermistor [4K-1.6K switch, 4K, 1.6K at 100K-200K]
17/05/2009 12:28:32	HFI: set Cernox thermistor [0.1-K nominal at 100K-200K]
18/05/2009 00:00:00	period: end of LEOP / start of Commissioning phase
01/06/2009 21:36:00	therm: end of reflectors anticontamination
02/06/2009 16:40:20	HFI: set the JFET PID heaters at 130K
02/06/2009 17:00:06	HFI: REU hardware test
02/06/2009 17:20:06	HFI: start the 4-K cooler compressors
02/06/2009 17:39:56	HFI: save REU configuration in EEPROM
03/06/2009 17:24:00	therm: SCS active cooldown starts
05/06/2009 17:28:18	slew: OCM-z (MAN-02a)
06/06/2009 18:19:21	HFI: set Cernox thermistor [4K-1.6K switch, 4K, 1.6K at 40K-100K]
09/06/2009 16:29:33	HFI: activate the 18K-4K heat switch nominal (repeated every day till 24/06)
09/06/2009 22:28:02	HFI: Single Event Upset
11/06/2009 21:15:00	HFI: set Cernox thermistor [0.1-K nominal at 40K-100K]
12/06/2009 03:00:00	HFI: QEC1 at 100K
12/06/2009 05:08:20	HFI: QEC1 at 100K [end]
12/06/2009 15:26:07	HFI: set Cernox thermistor [4K-1.6K switch, 4K, 1.6K at 19K-40K]
12/06/2009 15:32:10	HFI: PID setup [4K-1.6K switch PID activated at max power]
12/06/2009 15:33:31	HFI: PID setup [4-K PID nominal at max power]
15/06/2009 03:00:00	HFI: LFER_A_test
15/06/2009 04:54:52	HFI: LFER_A_test [end]
17/06/2009 16:46:58	slew: OCM-z touch-up (MAN-02b)
18/06/2009 18:29:57	HFI: LFER_A
19/06/2009 13:12:53	HFI: LFER_A [end]
22/06/2009 10:11:08	HFI: set Cernox thermistor [0.1-K nominal at 19K-40K]
23/06/2009 11:00:34	HFI: set Cernox thermistor [4K-1.6K switch, 4K, 1.6K at 9K-19K]
24/06/2009 10:40:15	HFI: PID setup [4K-1.6K switch PID deactivated]
24/06/2009 10:40:35	HFI: PID setup [4-K PID nominal]
24/06/2009 10:44:36	HFI: activate the 18K-4K heat switch nominal for LFI test [1st attempt]
24/06/2009 10:49:01	HFI: set the 4-K cooler amplitude demand at 3500 microns
24/06/2009 10:52:48	HFI: set the dilution cooler flow to FNOM2
24/06/2009 11:18:58	HFI: activate the 18K-4K heat switch nominal for LFI test [5th and last attempt]
26/06/2009 23:00:00	HFI: deactivate the 18K-4K heat switch nominal
27/06/2009 13:34:34	HFI: set Cernox thermistor [0.1K, 1.6K, 4K and 4K-1.6K switch at <9K]
28/06/2009 12:22:14	HFI: channel setup [R & C]
28/06/2009 12:53:22	HFI: 4-K cooler amplitude tuning 1/2
29/06/2009 09:38:15	HFI: 4-K cooler amplitude tuning 2/2
29/06/2009 11:58:41	HFI: channel setup [all 1.6-K thermometers]
29/06/2009 12:08:27	HFI: VI measurement [1.6-K thermometers]
29/06/2009 12:21:10	HFI: set the channel compression parameters [4-K thermometers]
29/06/2009 12:29:51	HFI: channel setup [all 4-K thermometers]
29/06/2009 12:38:30	HFI: VI measurement [4-K thermometers]
29/06/2009 13:28:40	HFI: channel setup [4-K PID nominal thermometer]
29/06/2009 13:31:24	HFI: PID setup [4-K PID nominal]
29/06/2009 13:41:00	HFI: channel setup [4-K PID nominal thermometer]
29/06/2009 14:56:15	HFI: channel setup [1.6-K PID nominal thermometer]
29/06/2009 14:58:58	HFI: PID setup [1.6-K PID nominal]
29/06/2009 15:07:42	HFI: set Cernox thermistor [0.1-K nominal at 3.5K-9K]
29/06/2009 15:45:02	HFI: PID setup [4-K PID nominal] (8 times hourly)
29/06/2009 21:59:57	HFI: LFER-B
30/06/2009 04:16:13	HFI: LFER-B [end]
30/06/2009 09:35:52	HFI: set Cernox thermistor [0.1-K nominal at 2.2K-3.5K]
30/06/2009 10:41:53	HFI: channel setup [1.6-K PID nominal thermometer]
30/06/2009 10:43:58	HFI: PID setup [1.6-K PID nominal]
30/06/2009 10:57:31	HFI: channel setup [0.1-K thermometers]
30/06/2009 12:35:07	HFI: set Cernox thermistor [0.1-K nominal at 1.4K-2.2K]
30/06/2009 13:42:12	HFI: PID setup [1.6-K PID nominal]
30/06/2009 14:20:51	HFI: VI measurement []
30/06/2009 18:09:59	HFI: set Cernox thermistor [0.1-K nominal at <1.4 K]
30/06/2009 18:20:51	HFI: VI measurement []
30/06/2009 19:10:01	HFI: QEC2 at 1K
30/06/2009 21:18:31	HFI: QEC2 at 1K [end]
01/07/2009 00:20:51	HFI: VI measurement
01/07/2009 04:20:51	HFI: VI measurement
01/07/2009 12:10:01	HFI: QEC3 at 300mK
01/07/2009 14:18:21	HFI: QEC3 at 300mK [end]
01/07/2009 14:18:31	HFI: channel setup [bolometers at 200mk]
01/07/2009 15:20:51	HFI: VI measurement []
01/07/2009 19:20:51	HFI: VI measurement []
01/07/2009 23:20:51	HFI: VI measurement []
02/07/2009 04:20:51	HFI: VI measurement []
02/07/2009 08:20:51	HFI: VI measurement []
02/07/2009 09:24:41	HFI: channel setup [0.1-K thermometers]
02/07/2009 10:35:26	HFI: channel setup [1.6-K PID nominal thermometer]
02/07/2009 10:36:58	HFI: PID setup [1.6-K PID nominal]
02/07/2009 11:15:46	slew: OCM Insertion (MAN-03a)
02/07/2009 12:34:36	HFI: set the dilution cooler flow to FNOM1
02/07/2009 13:58:16	HFI: set the channel compression parameters
03/07/2009 09:37:20	HFI: VI measurement [100-mK bolometers]
03/07/2009 10:10:01	HFI: channel setup [bolometers at 100mk]
03/07/2009 10:17:56	HFI: set the dilution cooler flow to FMIN2
03/07/2009 12:50:00	HFI: PID setup [bolo PID activated]
03/07/2009 13:09:45	HFI: channel setup [thermo PID1R]
03/07/2009 13:10:58	HFI: PID setup [dilu PID activated]
03/07/2009 13:26:04	HFI: channel setup [1.6-K PID nominal thermometer to 1.39 K]
03/07/2009 13:52:29	HFI: channel setup [1.6-K PID nominal thermometer to 1.41 K]
03/07/2009 14:03:43	HFI: set the channel compression parameters
03/07/2009 16:15:01	HFI: QEC4 at 100mK
03/07/2009 18:23:21	HFI: QEC4 at 100mK [end]
03/07/2009 19:05:51	HFI: VI measurement []
03/07/2009 20:05:26	HFI: PID setup [1.6-K PID nominal]
03/07/2009 22:00:02	HFI: pre-RAW72
03/07/2009 23:13:08	HFI: pre-RAW72 [end]
04/07/2009 02:00:00	therm: aphelion
04/07/2009 09:43:13	HFI: channel setup [1.6-K PID nominal thermometer to 1.39 K]
04/07/2009 09:45:27	HFI: PID setup [1.6-K PID nominal]
04/07/2009 20:26:32	HFI: channel setup [1.6-K PID nominal thermometer]
04/07/2009 22:30:01	HFI: CPVA_01
05/07/2009 01:55:49	HFI: CPVA_01 [end]
05/07/2009 10:47:08	HFI: channel setup [1.6-K PID nominal thermometer to 1.35 K]
05/07/2009 11:28:57	HFI: channel setup [1.6-K PID nominal thermometer to 1.34 K]
05/07/2009 13:25:03	HFI: IBTU_01
06/07/2009 09:47:49	HFI: IBTU_01 [end]
06/07/2009 09:58:51	HFI: channel setup [1.6-K PID nominal thermometer to 1.33 K]
06/07/2009 12:38:33	HFI: channel setup [1.6-K PID nominal thermometer to 1.35 K]
06/07/2009 13:15:21	HFI: channel setup [100-mK PID redundant to 98.5 mK]
07/07/2009 11:30:14	HFI: VI measurement [resistor, bolometers]
07/07/2009 13:24:37	HFI: set the channel compression parameters
07/07/2009 14:20:03	HFI: IBTU_02
08/07/2009 10:42:49	HFI: IBTU_02 [end]
08/07/2009 11:04:39	HFI: channel setup [1.6-K PID nominal thermometer to 1.39 K]
08/07/2009 11:16:02	HFI: PID setup [1.6-K PID redundant at 200 microW]
08/07/2009 11:24:40	HFI: channel setup [100-mK PID redundant to 100.0 mK]
08/07/2009 23:00:01	HFI: LFER-B
09/07/2009 17:39:59	HFI: LFER-B [end]
09/07/2009 19:30:01	HFI: IBTU_03
10/07/2009 00:00:00	Start of CPV phase
10/07/2009 15:08:03	HFI: IBTU_03 [end]
10/07/2009 16:00:01	HFI: PHTU_01
11/07/2009 00:01:47	HFI: PHTU_01 [end]
11/07/2009 10:24:00	HFI: channel setup [4-K PID nominal thermometer at 4.73 K]
11/07/2009 10:26:00	HFI: PID setup [deactivation of 4-K PID redundant] 
11/07/2009 10:50:16	HFI: 4KFT (bis)
11/07/2009 12:49:50	HFI: 4KFT (bis) [end]
11/07/2009 12:59:35	HFI: channel setup [1.6-K PID nominal thermometer to 1.37 K]
11/07/2009 13:05:25	HFI: PID setup [1.6-K PID redundant deactivated]
11/07/2009 14:10:01	HFI: XTLK_01
12/07/2009 00:34:38	HFI: Single Event Upset
12/07/2009 00:42:05	HFI: XTLK_01 [end]
12/07/2009 01:00:02	HFI: RAW72_02
12/07/2009 07:00:10	HFI: RAW72_02 [end]
12/07/2009 11:53:22	HFI: DPU soft reboot
12/07/2009 12:18:13	HFI: back to nominal
12/07/2009 13:02:33	HFI: channel setup [capacitor]
12/07/2009 13:22:13	HFI: set the channel compression parameters
12/07/2009 13:31:21	HFI: 01TO_01
13/07/2009 01:30:01	HFI: 01TO_01 [end]
13/07/2009 11:12:32	HFI: channel setup [4-K PID nominal thermometer at 4.80 K]
13/07/2009 13:03:50	HFI: channel setup [4-K PID nominal thermometer at 4.66 K]
13/07/2009 13:26:23	HFI: channel setup [4-K PID nominal thermometer at 4.64 K]
14/07/2009 09:50:15	HFI: PID setup [1.6-K PID redundant to 200 microW]
14/07/2009 13:12:01	HFI: PID setup [1.6-K PID redundant to 260 microW]
14/07/2009 13:14:22	HFI: channel setup [1.6-K PID nominal thermometer to 1.38 K]
15/07/2009 09:16:02	slew: OCM Insertion touch-up (MAN-03b)
15/07/2009 09:46:54	HFI: channel setup [4-K PID nominal thermometer at 4.66 K]
16/07/2009 09:21:29	HFI: PID setup [4-K PID redundant to 0.8 mW]
16/07/2009 09:23:30	HFI: channel setup [4-K PID nominal thermometer at 4.68 K]
17/07/2009 09:26:56	HFI: PID setup [4-K PID redundant to 1.2 mW]
17/07/2009 09:30:08	HFI: channel setup [4-K PID nominal thermometer at 4.70 K]
17/07/2009 11:12:09	HFI: channel setup [1.6-K PID nominal thermometer to 1.39 K]
17/07/2009 11:23:02	HFI: channel setup [thermometers]
17/07/2009 14:00:01	HFI: LFER_02
18/07/2009 08:40:59	HFI: LFER_02 [end]
18/07/2009 09:51:16	HFI: channel setup [1.6-K PID nominal thermometer to 1.40 K]
18/07/2009 19:10:19	therm: elephant
19/07/2009 11:09:09	HFI: channel setup [100-mK PID redundant to 100.5 mK]
20/07/2009 10:59:16	HFI: PID setup [1.6-K PID redundant deactivated]
20/07/2009 12:53:40	HFI: channel setup [1.6-K PID nominal thermometer to 1.37 K]
21/07/2009 10:05:59	HFI: channel setup [4-K PID nominal thermometer at 4.72 K]
21/07/2009 13:02:01	HFI: channel setup [bolo PID at 102.3 mK]
21/07/2009 13:11:48	HFI: PID setup [bolo PID]
22/07/2009 10:21:20	HFI: channel setup [1.6-K PID nominal thermometer to 1.38 K]
22/07/2009 10:26:49	HFI: PID setup [bolo PID]
23/07/2009 09:23:42	HFI: decrease the 4-K cooler amplitude demand to 3450 microns
23/07/2009 09:47:32	HFI: PID setup [4-K PID redundant at max power]
23/07/2009 09:54:04	HFI: channel setup [4-K PID nominal thermometer at 4.77 K]
23/07/2009 10:53:39	HFI: channel setup [4-K PID nominal thermometer at 4.80 K]
24/07/2009 10:44:08	HFI: PID setup [bolo PID]
24/07/2009 22:30:00	HFI: 4KTO
25/07/2009 08:30:00	HFI: 4KTO [end]
27/07/2009 01:56:42	HFI: Single Event Upset
27/07/2009 09:21:28	HFI: PID setup [4-K PID redundant to 0,84 mW]
27/07/2009 09:27:54	HFI: 4K memory HCT dump
27/07/2009 09:56:06	HFI: channel setup [bolo PID at 103.0 mK]
27/07/2009 13:01:32	HFI: set the channel compression parameters (increase of the HFI telemetry rate)
28/07/2009 09:31:34	HFI: channel setup [nominal bolo PID]
28/07/2009 09:38:16	HFI: channel setup [100-mK PID redundant to 101.0 mK]
28/07/2009 12:30:34	HFI: channel setup [nominal bolo PID]
28/07/2009 12:46:13	HFI: channel setup [nominal bolo PID]
31/07/2009 16:00:00	HFI: SUNI_1_1
01/08/2009 00:00:00	HFI: SUNI_1_2
01/08/2009 10:55:00	HFI: SPIN day 1 : RPM@1.05
02/08/2009 12:20:00	HFI: SPIN day 2 : RPM@0.95
03/08/2009 11:53:37	HFI: SPIN day 3 : RPM@1.00
03/08/2009 11:53:37	HFI: SPIN day 4 : RPM@1.00
05/08/2009 09:36:01	slew: 1.7' slews test
06/08/2009 10:00:00	HFI: SUNI_2_1
06/08/2009 17:31:31	HFI: 4-K Charge Regulator Unit anomalous shutdown
06/08/2009 18:00:00	HFI: SUNI_2_2
07/08/2009 12:20:01	HFI: set Cernox thermistor [4K-1.6K switch, 4K, 1.6K]
07/08/2009 13:40:43	HFI: 4-K CDE software reset
07/08/2009 13:45:43	HFI: activate the 4-K VCS
07/08/2009 13:46:23	HFI: select the harmonic forces
07/08/2009 13:47:56	HFI: set the 4-K cooler amplitude demand at 2000 microns
07/08/2009 13:49:28	HFI: start the 4-K cooler compressors
07/08/2009 13:53:23	HFI: set the 4-K cooler amplitude demand at 2500 microns
07/08/2009 13:55:16	HFI: set the 4-K cooler amplitude demand at 3000 microns
07/08/2009 13:58:50	HFI: 4-K getter on
07/08/2009 17:54:56	HFI: stop the 4-K cooler compressors
07/08/2009 17:57:22	HFI: deactivate the 4-K VCS
07/08/2009 17:58:17	HFI: set the 4-K cooler amplitude demand at 3500 microns
07/08/2009 17:59:41	HFI: start the 4-K cooler compressors
07/08/2009 18:41:27	HFI: PID setup [4-K PID redundant]
08/08/2009 10:05:25	HFI: stop the 4-K cooler compressors
08/08/2009 10:07:41	HFI: activate the 4-K VCS
08/08/2009 10:08:09	HFI: select the harmonic forces
08/08/2009 10:08:52	HFI: set the 4-K cooler amplitude demand at 3450 microns
08/08/2009 10:10:26	HFI: start the 4-K cooler compressors
08/08/2009 10:48:56	HFI: set Cernox thermistor [4K-1.6K switch, 4K, 1.6K]
08/08/2009 11:48:33	HFI: PID setup [dilu PID activated]
08/08/2009 11:55:25	HFI: PID setup [4-K PID redundant to 0,84 mW]
08/08/2009 14:05:36	period: first light survey
09/08/2009 03:58:53	star tracker switchover #1
11/08/2009 09:44:16	HFI: channel setup [100-mK thermometers]
12/08/2009 10:24:46	HFI: channel setup [100-mK PID redundant]
12/08/2009 10:29:52	HFI: PID setup [dilu PID]
12/08/2009 14:13:45	<span style="background:#00FF00">period: DPC survey#1 (ring#240)</span>
14/08/2009 12:30:10	slew: orbit maintenance manoeuvre #1
18/08/2009 11:43:42	long ring #440 for CDMU patch day 1
18/08/2009 12:39:09	sat: CDMU patch (day 1)
18/08/2009 12:39:20	sat: CDMU patch (end of day 1)
19/08/2009 11:49:40	long ring #474 for CDMU patch day 2
19/08/2009 12:45:31	sat: CDMU patch (day 2)
19/08/2009 12:45:34	sat: CDMU patch (end of day 2)
20/08/2009 11:53:46	long ring #509 for CDMU patch day 3
20/08/2009 12:45:32	sat: CDMU patch (day 3)
20/08/2009 12:45:36	sat: CDMU patch (end of day 3)
21/08/2009 11:56:03	long ring #544 for CDMU patch
21/08/2009 14:18:48	HFI: clock resynchronisation
22/08/2009 09:36:51	dtcp: transition to 3 h DTCP duration
31/08/2009 12:09:14	long ring #897 to allow the execution of the ACMS warm reset
31/08/2009 13:40:00	thrusters catbed heaters switched off by error
31/08/2009 13:40:01	sat: FOG telemetry disabled by error
31/08/2009 13:41:05	long ring #898 to allow the execution of the ACMS warm reset
04/09/2009 15:01:00	thruster under/over performed
11/09/2009 21:00:41	slew: orbit maintenance manoeuvre #2
16/09/2009 20:31:41	thrusters catbed heaters switched back on
18/09/2009 19:06:15	planet: HFI Crab scan#1
20/09/2009 19:36:00	dtcp: extra pass CEB
20/10/2009 23:44:24	therm: HFI elephant LONG deltaT>10uK
22/10/2009 19:34:56	planet: HFI Mars scan#1
24/10/2009 11:01:39	therm: HFI elephant LONG deltaT>10uK
25/10/2009 15:30:49	planet: HFI Jupiter scan#1
28/10/2009 23:06:00	thruster under/over performed
01/11/2009 00:41:04	planet: HFI Neptune scan# 1
20/11/2009 18:52:25	long ring #3589 (29 hours) : no slew due to MTL uplink problem
21/11/2009 23:39:54	slew : resumption of SL after the MTL problem
22/11/2009 19:48:41	slew to SGR position
23/11/2009 08:47:54	slew to catchup
23/11/2009 19:48:10	SGR (end): back to nominal scanning law
25/11/2009 05:56:00	HFI: 1.6-K stage event
26/11/2009 14:32:00	thruster under/over performed
04/12/2009 17:49:29	slew: orbit maintenance manoeuvre #3
05/12/2009 09:43:53	planet: HFI Uranus scan# 1
16/12/2009 20:42:11	sat: FOG telemetry enabled back
21/12/2009 19:59:54	HFI: set the channel compression parameters
22/12/2009 11:46:18	dtcp: extra pass
03/01/2010 00:00:00	therm: perihelia
04/01/2010 00:25:49	planet: HFI Saturn scan#1
11/01/2010 19:16:32	star tracker switchover #2
15/01/2010 21:57:45	slew: orbit maintenance manoeuvre #4
25/01/2010 21:16:41	sat: transponder always on
08/02/2010 20:51:32	<span style="background:#00FF00">period: DPC survey#2 (ring#5721)</span>
26/02/2010 12:43:37	star tracker switchover #3
26/02/2010 23:02:49	slew: orbit maintenance manoeuvre #5
06/03/2010 05:04:43	planet: HFI Crab scan#2
01/04/2010 06:41:00	thruster under/over performed
02/04/2010 13:48:21	therm: major elephant
05/04/2010 10:00:00	solar flare: geomagnetic storm
10/04/2010 10:36:35	planet: HFI Mars scan#2
26/04/2010 17:55:37	slew: orbit maintenance manoeuvre #6
29/04/2010 17:34:58	therm: SCS TSA setpoint increase
06/05/2010 17:02:08	therm: SCS TSA setpoint increase
13/05/2010 16:30:13	therm: SCS TSA setpoint increase
17/05/2010 17:51:09	planet: HFI Neptune scan# 2
12/06/2010 01:00:00	solar flare
12/06/2010 01:57:16	planet: HFI Saturn scan#2
24/06/2010 18:31:09	therm: SCS TSA setpoint increase
30/06/2010 19:00:53	planet: HFI Uranus scan#2
03/07/2010 20:01:43	planet: HFI Jupiter scan#2
06/07/2010 11:00:00	therm: aphelion
14/07/2010 20:45:50	dtcp: extra pass
15/07/2010 18:13:43	therm: SCS TSA setpoint increase
18/07/2010 04:28:00	thruster under/over performed
21/07/2010 21:08:44	therm: moon transit
22/07/2010 00:42:59	therm: moon transit (end)
24/07/2010 16:36:11	therm: HFI elephant LONG deltaT>10uK
29/07/2010 17:45:52	therm: SCS TSA setpoint increase
03/08/2010 19:00:00	solar flare: possible geomagnetic storm
11/08/2010 10:14:37	long ring #11149 for SCS switchover
11/08/2010 11:00:00	therm: SCS switchover
11/08/2010 11:00:30	HFI: PID setup [1.6-K PID nominal and 4-K PIDs deactivated]
11/08/2010 17:17:50	HFI: deselect all 4-K harmonics for VCS 
11/08/2010 18:50:15	HFI: select all 4-K harmonics for VCS 
11/08/2010 18:52:44	HFI: PID setup [1.6-K PID nominal and 4-K PIDs reactivated]
11/08/2010 21:34:10	dtcp: extra pass CEB
12/08/2010 00:06:59	HFI: PID setup [4-K PID redundant set to 1.14 mW]
12/08/2010 19:27:28	<span style="background:#00FF00">period: DPC survey#3 (ring#11195)</span>
13/08/2010 23:59:21	slew to interleaved pointings for survey#3
14/08/2010 11:00:00	solar flare
16/08/2010 18:00:09	slew: orbit maintenance manoeuvre #7
18/08/2010 08:35:00	solar flare
25/08/2010 17:51:58	HFI: 4-K memory HCT dump
18/09/2010 18:28:04	planet: HFI Crab scan#3
04/10/2010 04:25:00	HFI: 1.6-K stage event
04/10/2010 04:57:05	therm: HFI elephant SHORT deltaT>10uK
06/10/2010 19:00:46	dtcp: extra pass CEB
11/10/2010 13:26:47	long ring  #13333: manoeuvre failed
03/11/2010 08:11:34	planet: HFI Neptune scan#3
04/11/2010 13:20:14	therm: change of the DPU2 thermal control
04/11/2010 19:19:00	dtcp: extra pass NNO
11/11/2010 12:45:00	therm: SCS TSA setpoint increase
21/11/2010 03:02:51	therm: HFI elephant LONG deltaT>10uK
24/11/2010 11:23:01	long ring  #14627 for star tracker tests
24/11/2010 12:29:29	star tracker switchover for intercalibration test (STR1->STR2)
24/11/2010 20:23:53	slew back
25/11/2010 17:02:45	long ring  #14653 for star tracker tests
25/11/2010 17:59:37	star tracker switchover STR2->STR1
26/11/2010 01:55:21	slew back to nominal scanning strategy
27/11/2010 20:20:00	solar flare
02/12/2010 13:21:13	slew: orbit maintenance manoeuvre #8
06/12/2010 20:38:00	planet: HFI Jupiter scan#3
10/12/2010 14:55:44	planet: HFI Uranus scan#3
15/12/2010 08:19:00	thruster under/over performed
15/12/2010 10:30:00	thruster under/over performed
22/12/2010 16:33:14	HFI: Single Event Upset
23/12/2010 15:44:53	therm: SCS TSA setpoint increase
03/01/2011 19:00:00	therm: perihelia
18/01/2011 05:25:24	planet: HFI Saturn scan#3
24/01/2011 13:20:09	therm: SCS warm radiator warming
24/01/2011 17:28:00	dtcp: extra pass CEB
28/01/2011 02:00:00	solar flare
31/01/2011 14:06:49	slew: orbit maintenance manoeuvre #9
01/02/2011 13:31:33	therm: SCS TSA setpoint increase
08/02/2011 20:55:48	<span style="background:#00FF00">period: DPC survey#4 (ring#16692)</span>
10/02/2011 02:48:02	HFI: 1.6-K stage event
15/02/2011 06:00:00	solar flare: X ray solar flare
18/02/2011 00:00:00	solar flare: X ray solar flare
24/02/2011 14:05:28	therm: SCS TSA setpoint increase
28/02/2011 23:20:41	slew: acceleration for Crab scan
06/03/2011 04:12:44	planet: HFI Crab scan#4
07/03/2011 21:00:00	solar flare: coronal mass ejection
10/03/2011 19:44:00	dtcp: extra pass CEB
15/03/2011 17:42:31	therm: SCS desorption power gap jump
15/03/2011 20:35:00	dtcp: extra pass
18/03/2011 15:15:00	dtcp: long pass (5h30mn)
21/03/2011 03:30:00	solar flare: radiation event
21/03/2011 21:10:17	slew: orbit maintenance manoeuvre #10
31/03/2011 21:02:52	therm: SCS TSA setpoint increase
07/04/2011 20:38:16	therm: SCS TSA setpoint increase
14/04/2011 21:03:33	therm: SCS TSA setpoint increase
21/04/2011 20:33:20	therm: SCS TSA setpoint increase
10/05/2011 16:48:13	therm: SCS TSA setpoint increase
20/05/2011 14:18:21	planet: HFI Neptune scan#4
04/06/2011 22:00:00	solar flare: geomagnetic storm
07/06/2011 07:00:00	solar flare: solar proton enhancement
23/06/2011 04:55:03	thruster under/over performed
29/06/2011 17:03:39	planet: HFI Saturn scan#4
04/07/2011 15:00:00	therm: aphelion
05/07/2011 22:00:17	planet: HFI Uranus scan#4
25/07/2011 17:38:00	slew: orbit maintenance manoeuvre #11
27/07/2011 16:51:00	on board mass properties update
29/07/2011 17:13:00	slew to survey#5
29/07/2011 17:43:32	<span style="background:#00FF00">period: DPC survey#5 (ring#21721)</span>
02/08/2011 07:00:00	solar flare: M1 class
03/08/2011 21:09:40	planet: HFI Jupiter scan#4
04/08/2011 04:30:00	solar flare
04/08/2011 15:55:38	therm: SCS TSA setpoint increase
08/08/2011 16:53:00	SGR: scanning law interrupted
08/08/2011 17:32:00	slew back
08/08/2011 18:00:00	solar flare: X7 class
08/08/2011 18:44:00	planet: HFI Jupiter scan#4 (bis)
09/08/2011 01:54:00	slew to the catchup position
09/08/2011 08:00:00	solar flare
09/08/2011 17:23:00	SGR (end): back to nominal scanning law
06/09/2011 02:20:00	solar flare: M5.3
06/09/2011 23:00:00	solar flare: X2.1
13/09/2011 09:34:22	planet: HFI Crab scan#5
17/09/2011 08:00:00	solar flare: magnetic storm
20/09/2011 09:43:00	slew setting-up the second crab scan 
20/09/2011 10:41:38	planet: HFI Crab scan#5 (bis)
21/09/2011 16:43:41	planet: HFI Crab scan#6
21/09/2011 18:33:00	dtcp: extra pass
22/09/2011 12:00:00	solar flare: complexe event
24/09/2011 00:42:34	slew back to survey 5 scanning strategy 
26/09/2011 12:29:45	slew : station keeping manoeuvre half failed
04/10/2011 12:35:07	slew: orbit maintenance manoeuvre #12 bis
11/10/2011 12:35:13	HFI:  temperature steps on bolometer PID
12/10/2011 10:30:00	HFI:  EOL-Fast VI at 100mK
22/10/2011 14:00:00	solar flare: small event
27/10/2011 12:38:38	therm: SCS TSA setpoint increase
04/11/2011 00:00:00	solar flare: X2 class
07/11/2011 09:00:00	HFI : pre-EOL-LFER-1
07/11/2011 11:06:54	HFI : pre-EOL-LFER-1 [end]
08/11/2011 10:42:00	HFI:  EOL-IBTU
08/11/2011 16:10:03	HFI:  EOL-IBTU [end]
14/11/2011 09:00:00	HFI:  2nd pre-EOL-LFER-1
14/11/2011 11:08:42	HFI:  2nd pre-EOL-LFER-1 [end]
14/11/2011 23:58:33	slew: beginning of dwell duration compression phase 
18/11/2011 00:00:00	HFI: EOL-LFER-1
20/11/2011 14:33:42	HFI: EOL-LFER-1 [end]
26/11/2011 07:00:00	solar flare: Radiation storm and Coronal Mass Ejection
28/11/2011 17:10:00	dtcp: extra pass
07/12/2011 17:00:00	spinup test at 1.4RPM
08/12/2011 10:00:00	planet: HFI Mars scan @ 1.4 RPM
09/12/2011 16:43:00	star tracker unhealthy
16/12/2011 18:35:00	slew: back to scanning law
17/12/2011 18:00:00	planet: HFI Mars scan#3
23/12/2011 15:10:06	HFI: He3 flow rate 0 -> 1
25/12/2011 14:45:16	HFI:  HFI: PID setup [1.6-K PID redundant add 153.4 uW]
25/12/2011 20:00:00	solar flare: C-class
27/12/2011 11:42:10	HFI: temperature steps on the bolometer plate II
27/12/2011 13:48:45	HFI: temperature steps on the bolometer plate II [end]
29/12/2011 14:24:29	HFI: He3 flow rate 1 -> 2
02/01/2012 13:58:00	HFI: He3 flow rate 2 -> 3
03/01/2012 14:20:19	HFI: temperature steps on bolo plate III (1st part)
04/01/2012 13:52:00	HFI: He3 flow rate 3 -> 4
04/01/2012 16:00:00	HFI: temperature steps on bolo plate III (2nd part)
06/01/2012 00:00:00	HFI: temperature steps on bolo plate III (2nd part) [end]
07/01/2012 14:10:52	HFI: He3 flow rate 4 -> 5
08/01/2012 07:20:00	slew to Jupiter deep annuli scan
09/01/2012 14:23:36	HFI: He3 flow rate 5 -> 6
10/01/2012 04:20:00	planet: HFI Jupiter scan
11/01/2012 14:48:04	HFI: He3 flow rate 6 -> 7
13/01/2012 15:04:33	HFI: HFI: PID setup [1.6-K PID and dilu PID deactivated]
14/01/2012 16:00:00	<span style="background:#00FF00">period: end of HFI nominal operations</span>
16/01/2012 15:00:00	HFI channel setup [100mK thermometers]
16/01/2012 15:00:12	HFI: bolo plate at 110mK
16/01/2012 15:45:09	HFI: bolo plate at 300mK
17/01/2012 14:13:54	HFI: set Cernox thermistor [0.1-K nominal at 1.4K]
18/01/2012 15:20:10	HFI: set the channel compression parameters
18/01/2012 16:15:11	HFI: EOL-Fast VI at 300 mK
19/01/2012 14:30:52	HFI: channel setup [bolometers]
20/01/2012 19:58:08	HFI: EOL-RAW54 at 300mK
21/01/2012 13:29:58	HFI: EOL-RAW54 at 300mK [end]
22/01/2012 00:18:00	therm: loss of satellite thermal control
22/01/2012 15:00:00	CDMU PM-A reset
22/01/2012 22:59:58	HFI: DPU soft reboot
23/01/2012 05:00:00	solar flare: major one 
25/01/2012 14:45:00	HFI: set the channel compression parameters
27/01/2012 19:00:00	solar flare
30/01/2012 21:53:00	period: slew to survey#6
07/02/2012 00:00:00	slew: large orbit correction manoeuvre
10/02/2012 16:30:00	HFI: EOL-Fast VI at 600 mK
14/02/2012 14:20:00	therm: CDMU PM-A switch over
20/02/2012 10:00:00	HFI: set the channel compression parameters (EOL#0)
20/02/2012 14:57:00	slew: touchup manoeuvre
21/02/2012 15:00:00	HFI: set the channel compression parameters (EOL#1)
21/02/2012 15:16:00	HFI: set the channel compression parameters (EOL#2)
22/02/2012 16:50:00	HFI: set the channel compression parameters (EOL#3)
23/02/2012 15:20:00	HFI: set the channel compression parameters (EOL#4)